Write a system of two equations in two unknowns for each
problem. Solve each system by the method of your choice
Girls and boys. There are 385 surfers in Surf City. Twothirds
of the boys are surfers and one-twelfth of the girls
are surfers. If there are two girls for every boy, then how
many boys and how many girls are there in Surf City?
Can someone please help me solve this

Answer by oberobic(2304)

B = number of boys
G = number of girls
S = number of surfers = 385
.
2/3B + 1/12G = S = 385
.
There are 2 girls for each boy.
2B = G .or.
B = 1/2G
.
Substituting into the total equation...
2/3B + 1/12G = 385
.
Multiplying through by 12 to remove the fractions...
8B + G = 4620
.
Substituting for B with 1/2G
8*1/2G + G = 4620
4G + G = 4620
5G = 4620
G = 924
.
Since we know there are twice as many girls as boys,
B = G/2 = 924/2 = 462
.
Now we can check to see if we substitute these value that we find there are 385 surfers...
.
2/3*B = 308
1/12*G = 77
Total=385